Iron forms two compounds with chlorine, compound X and compound Y. When 3.00 g of compound X was analyzed it was found that 1.68 g of CI combined with 1.72 g of Fe. When 5.00 g of compound Y was analyzed it was found 3.28 g CI combined with 1.72 g of Fe. How many g of CI would combine with 1.00 g of Fe? Do for both X and Y using law of multiple proportion
1
Expert's answer
2020-09-10T05:59:20-0400

Compound X

1.68 Cl combines with 1.72 Fe

a Cl combines with 1.00 Fe

a = 1.68/1.72 = 0.977 (g) of Cl


Compound Y

3.28 Cl combines with 1.72 Fe

b Cl combines with 1.00 Fe

b = 3.28/1.72 = 1.91 (g) of Cl


a : b = 1 : 2